OpenStudy (anonymous):

Solve A = 1/2p + t for p in terms of A and t.

OpenStudy (anonymous):

p = 2(A – t) p = 2A – t p = A – 2t p = \[\frac{ A-t }{ 2}\]

OpenStudy (anonymous):

Is it B?

OpenStudy (anonymous):

Try to get P all by itself on one side of the equation. try it, step by step.

OpenStudy (anonymous):

Is it A?

OpenStudy (anonymous):

Is that what you got when you solved for p?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

There you go. It's A. Good job :)

OpenStudy (anonymous):

Thank you (: